OpenStudy (anonymous):

Susan has watered fifteen of her twenty houseplants. What is fifteen-twentieths in simplest form?

OpenStudy (anonymous):

Since Susan watered 15 of her 20 houseplants, 20 - 15 = 5 plants remain. Fifteen twentieths in simplest form is 15/20 : 5 = 3/4. So, in the simplest form, fifteen twentieths are three fourths. (3/4)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can I get a medal?

OpenStudy (anonymous):

your welcome :)